# Steering the Herd: A Framework for LLM-based Control of Social Learning

### v3

@@ Abstract
Algorithms increasingly serve as information mediators – from social media feeds and targeted advertising to the increasing ubiquity of LLMs. This engenders a joint process where agents combine private, algorithmically-mediated signals with observational learning from peers to arrive at decisions. To study such settings, we introduce a model of controlled sequential social learning in which an information-mediating planner (e.g., an LLM) controls the information structure of agents while they also learn from the decisions of earlier agents. The planner may seek to improve social welfare (an altruistic planner) or to induce a specific action the planner prefers (a biased planner). Our framework presents a new optimization problem for social learning that combines dynamic programming with decentralized action choices and Bayesian belief updates.
In this setting, we prove the convexity of the value function and characterize the optimal policies of altruistic and biased planners, which attain desired tradeoffs between the costs they incur and the payoffs they earn from induced agent choices. The characterization reveals that the optimal planner operates in different modes depending on the range of belief values. The modes include investing the maximum allowed resource, not investing any resource, or the investment increasing or decreasing with increase in the belief. Notably, for some ranges of belief the biased planner even intentionally obfuscates the agents’ signals. Even under stringent transparency constraints—information parity with individuals, no lying or cherry‑picking, and full observability—we show that information mediation can substantially shift social welfare in either direction. We complement our theory with simulations in which LLMs act as both planner and agents. Notably, the LLM-based planner in our simulations exhibits emergent strategic behavior in steering public opinion that broadly mirrors the trends predicted, though key deviations suggest the influence of non-Bayesian reasoning—consistent with the cognitive patterns of both human users and LLMs trained on human-like data. Together, we establish our framework as a tractable basis for studying the impact and regulation of LLM information mediators that corresponds to real behavior.
@@ 1 Introduction
Large Language Models (LLMs) are increasingly deployed as information mediators in socially-critical functions, from search engines and news aggregators (Zhang et al., 2024a; Gao et al., 2024) to personal assistants offering medical or political advice (Huo et al., 2025; Haupt & Marks, 2023; Schiele et al., 2024; Argyle et al., 2023). While promising, LLMs have also demonstrated the ability to be persuasive (Potter et al., 2024; Carrasco-Farre, 2024; Rogiers et al., 2024), manipulative (Williams et al., 2025; Liu et al., 2025; Jones & Bergen, 2024), and strategic (Lorè & Heydari, 2024; Payne & Alloui-Cros, 2025; Zhang et al., 2024b) in their user interactions. When deployed at scale, this algorithmic influence does not occur in a vacuum. It intersects with the organic spread of information as people observe and learn from one another, creating complex social dynamics that can magnify or mitigate the models’ impact.
@@
Crucially, however, a user’s decision is not based solely on this private signal. They also observe the actions of those who came before them—if the last ten users decided to patronize the restaurant, the next user will infer that they likely received positive information. This phenomenon, known as social learning, is a powerful force in shaping human opinion in a variety of settings including the adoption of vaccines (e.g., Rao et al. (2007); Bauch & Bhattacharyya (2012)), new technology (e.g. Gillingham & Bollinger (2021); Weber (2012)), and the evolution of political and moral opinions (e.g., Brady et al. (2021); Guilbeault et al. (2018)). Our work’s central contribution is to study the strategic interaction between a planner who controls private information and a population of agents who engage in social learning. The planner must therefore anticipate how its actions will not only influence the current user but also create an information externality that shapes the public belief confronting all future users.
@@ 1.1 Contributions and Outline
    Empirical Validation and Strategic Analysis Using LLMs. We conduct simulations where LLMs act as both planner and agents. Our experiments demonstrate three key findings: (a) A planner that accounts for social learning can dramatically influence public opinion and social welfare, far more than a myopic one. (b) The strategic behavior that emerges from the LLM planner largely aligns with our theoretical predictions, suggesting the model is robust to more realistic, non-Bayesian agent behavior. (c) LLMs exhibit sophisticated strategic reasoning, highlighting both their potential and the societal risks they present.
@@ 2 Related Work
We build on three main fields: social learning, information design, and some nascent work on LLMs.
Social Learning
@@
Bikhchandani et al. (1992) and Banerjee (1992) introduced the sequential social learning framework in economics, and subsequent work has thoroughly characterized social-learning dynamics in more general settings (e.g., Smith & Sørensen (2000); Arieli & Mueller-Frank (2021)). This literature has also studied many variations in the kind of social information available to agents and various biases and misperceptions of the agents (see Dasaratha & He (2022); Bistritz et al. (2022); Eyster & Rabin (2010) as well as Bikhchandani et al. (2024) and the references therein).
@@
The literature on the control of social learning is relatively limited (Wei & Anastasopoulos, 2022; Krishnamurthy, 2012; Bhatt & Krishnamurthy, 2021). Perhaps the closest work to ours is Wei & Anastasopoulos (2022) which takes a mechanism design approach that assumes two-way communication between the planner and agents (i.e., agents can report their private signals to the planner, and the planner can provide action recommendations based on all previous agents’ reports). Our model does not rely on two-way communication between the agents and the planner, making it more suited for information-mediating algorithms which are often invisible or black-boxed to the users.
Information Design
@@
In our setting, the planner optimizes their objective function by choosing the agents’ private signal precision. This is a constrained form of more classical information design questions (Bergemann & Morris, 2019). These models came into prominence with Kamenica & Gentzkow (2011)’s model of Bayesian persuasion — a planner chooses the signal structure informing a rational, Bayesian agent about an unknown state. Our planner’s problem is then a costly, constrained information design problem in each period, where the planner must choose a binary, symmetric signal at some cost. While many variations of information design have been studied (e.g. Castiglioni et al. (2020); Agrawal et al. (2023); Celli et al. (2020)), our primary innovation is to incorporate social learning between the agents.
LLMs
@@
More recently, some works have specifically studied LLMs in the context of information design (Harris et al., 2025; Dutting et al., 2025). This builds on the notion that LLMs can mimic humans as economic agents (Dillion et al., 2023; Horton, 2023). The most similar work in this area is Dutting et al. (2025). They study information design where a sender chooses both a signal structure and a ’framing’ that conditions the posterior belief of the receiver, possibly non-Bayesian. They implement an LLM-based approach for the computationally difficult task of optimizing over the framing space. Their work does not address our main considerations of repeated interactions or social-learning; but we adapt parts of their methodology for our empirical study in Section 6.
@@ 3.1 Agents’ Decision Problems
Informed by $\mathcal{H}_{i}$ (equivalently, by the Markov property, $b_{i}$) and $q_{i}$, agent $i$ chooses action $a_{i}\in\Omega$ so as to maximize her utility. If both actions fetch the same utility, she chooses the action that matches $s_{i}$. Agent $i$ obtains a private belief $\tilde{b}_{i}$ about $\omega$ (see Appendix B.1) using $\mathcal{H}_{i}$ and her private signal $s_{i}$ of precision $q_{i}$. Using this private belief, she chooses the action corresponding to the state of the world that is more likely as per her posterior belief $\tilde{b}_{i}$. In other words, she chooses $a_{i}=G$ if $\tilde{b}_{i}>0.5$; $a_{i}=B$ if $\tilde{b}_{i}<0.5$; and $a_{i}=s_{i}$ if $\tilde{b}_{i}=0.5$. Substituting for $\tilde{b}_{i}$, we find:
@@ 3.2 The Planner’s Problem
The optimal utility and policy of the altruistic planner are then defined as the supremum and arg supremum of $V_{A}^{\pi}(\cdot)$ over $\pi\in\Pi$ where $\Pi$ is the set of all policies.
@@
The difference between the biased planner’s problem and the altruistic planner’s is in their objectives and, therefore, in the cost and reward functions. Refer to the example for a political campaign in Section A for elucidation of a biased planner. The biased planner seeks to induce action $G$ from each agent regardless of $\omega.$ When an agent chooses action $G$ (respectively, $B$), the planner, incurs cost 0 (respectively, $C>0$), regardless of $\omega$. The biased planner can make a private signal more or less precise than the baseline value of $p$, both of which incur costs.Any choice of precision other than $p$ incurs a cost for the biased planner as it requires him to tailor the ad to an agent, which in turn needs research on how the agent best understands any content. The biased planner incurs cost $\beta(|q_{i}-p|)$ for choosing signal precision $q_{i}$, with $\beta(0)=0$.
@@
There are a few noteworthy assumptions we make in our model: (1) The planner has the exact history as the agents. This is limiting when the planner could have richer data, but not restrictive when all parties rely on the same public signals, such as early findings or survey data. (2) The planner can change agents’ signal precision, but signals are still passed through a binary symmetric channel. This is limiting if planners can cherry-pick, censor, or falsify signals, but not restrictive when their role is only to improve the quality of noisy data sources without altering content. (3) The planner’s control choices are fully observable. This is limiting when covert policies or framing are possible, but not restrictive in more transparent environments.
@@ 4 Optimal Altruistic Policies
The proof of Theorem 2 (Appendix C.3) is quite involved and may be of independent interest. The challenge is rooted in the dependence of agents’ actions on the public belief. In contrast, if the actions did not depend on the public belief process (e.g., as in Nyarko (1994)), the expected utility is linear function of the belief state, and the convexity of the value function then directly follows. The convexity of the value function is instrumental in characterizing the optimal policy.
@@
Furthermore, if $t_{M}<0.5$, then $d_{A}<t_{M}$. Proof in Appendix C.5.
@@ 5 Optimal Biased Policies
There exist $t_{1},\dots,t_{5}\in(0,p)$ with $t_{1}<1-p\leq t_{2}\leq t_{3}<0.5\leq t_{4}\leq t_{5}<p$ so that:
@@
Theorem 5 (proven in Appendix C.8) shows five possible phases.
@@ 6 Evaluation via LLM-based Simulation
To understand the implications and applicability of our theoretical results, we return to our motivating case: LLMs. We now utilize LLMs to empirically study the policies of information-mediating planners and the resultant dynamics in a setting with more realistic, non-Bayesian agents. Our evaluation proceeds in three steps: we first analyze the behavior of LLM agents to identify key deviations from Bayesian rationality, then examine how an LLM planner adapts its strategy in response, and finally evaluate the resulting impact on social welfare.
To simulate the controlled social learning setting, we operationalize our theoretical model using a scenario where agents decide whether to buy a new car. The abstract concept of signal precision ($q_{i}$) is translated into a concrete, context-aware message. We use LLMs in three distinct roles (figure 1(a)): (1) the agents, who make decisions based on public history and private messages; (2) the planner, who chooses a precision level based on the history; and (3) a belief oracle. Given the planner’s chosen precision $q_{i}$ and the agent’s profile, the oracle generates a tailored message (a textual car review) designed to have that specific level of informativeness.
We compare the outcomes of these LLM-based simulations with numerical evaluations of the planner’s MDP. In both LLM and numerical experiments, we assume a linear cost function, $\beta(q)=k|q-p|$, and vary $k$, baseline precision $p$, and discount factor $\delta$, with the cost of an incorrect action $C$ fixed at $1$. See Appendix E for further detail on experimental setup and prompting.
Figure 1: (a) A system diagram for our LLM simulations. The instance parameters detail are color-coded according to which of the LLM roles uses them. (b) Here we show the change in LLM (solid) and Bayesian (dashed) agents’ beliefs after receiving a positive (black) or negative (red) signal.

### v4

@@ Abstract
Algorithms increasingly serve as information mediators – from social media feeds and targeted advertising to the increasing ubiquity of LLMs. This engenders a joint process where agents combine private, algorithmically-mediated signals with observational learning from peers to arrive at decisions. To study such settings, we introduce a model of controlled sequential social learning in which an information-mediating planner (e.g., an LLM) controls the information structure of agents while they also learn from the decisions of earlier agents. The planner may seek to improve social welfare (an altruistic planner) or to induce a specific action the planner prefers (a biased planner). Our framework presents a new optimization problem for social learning that combines dynamic programming with decentralized action choices and Bayesian belief updates. In this setting, we prove the convexity of the value function and characterize the optimal policies of altruistic and biased planners, which attain desired tradeoffs between the costs they incur and the payoffs they earn from induced agent choices. The characterization reveals that the optimal planner operates in different modes depending on the range of belief values. The modes include investing the maximum allowed resource, not investing any resource, or the investment increasing or decreasing with increase in the belief. Notably, for some ranges of belief the biased planner even intentionally obfuscates the agents’ signals. Even under stringent transparency constraints—information parity with individuals, no lying or cherry‑picking, and full observability—we show that information mediation can substantially shift social welfare in either direction. We complement our theory with simulations in which LLMs act as both planner and agents. Notably, the LLM-based planner in our simulations exhibits emergent strategic behavior in steering public opinion that broadly mirrors the trends predicted, though key deviations suggest the influence of non-Bayesian reasoning—consistent with the cognitive patterns of both human users and LLMs trained on human-like data. Together, we establish our framework as a tractable basis for studying the impact and regulation of LLM information mediators that corresponds to real behavior.
@@ 1 Introduction
Large Language Models (LLMs) are increasingly deployed as information mediators in socially-critical functions, from search engines and news aggregators (Zhang et al., 2024a; Gao et al., 2024) to personal assistants offering medical or political advice (Huo et al., 2025; Haupt and Marks, 2023; Schiele et al., 2024; Argyle et al., 2023). While promising, LLMs have also demonstrated the ability to be persuasive (Potter et al., 2024; Carrasco-Farre, 2024; Rogiers et al., 2024), manipulative (Williams et al., 2025; Liu et al., 2025; Jones and Bergen, 2024), and strategic (Lorè and Heydari, 2024; Payne and Alloui-Cros, 2025; Zhang et al., 2024b) in their user interactions. When deployed at scale, this algorithmic influence does not occur in a vacuum. It intersects with the organic spread of information as people observe and learn from one another, creating complex social dynamics that can magnify or mitigate the models’ impact.
@@
Crucially, however, a user’s decision is not based solely on this private signal. They also observe the actions of their predecessors—if the last ten users decided to patronize the restaurant, the next user will infer that they likely received positive information. This phenomenon, known as social learning, is a powerful force in shaping human opinion in a variety of settings including the adoption of vaccines (e.g., Rao et al. (2007); Bauch and Bhattacharyya (2012)), new technology (e.g. Gillingham and Bollinger (2021); Weber (2012)), and political and moral opinions (e.g., Brady et al. (2021); Guilbeault et al. (2018)). Our work’s central contribution is to study the strategic interaction between a planner who controls private information and a population of agents who engage in social learning. The planner must therefore anticipate how its actions will not only influence the current user but also create an information externality that shapes the public belief confronting all future users.
@@
    Empirical Validation and Strategic Analysis Using LLMs. We conduct simulations where LLMs act as both planner and agents. Our experiments demonstrate three key findings: (a) A planner that accounts for social learning can dramatically influence public opinion and social welfare, far more than a myopic one. (b) The strategic behavior that emerges from the LLM planner largely aligns with our theoretical predictions, suggesting the model is robust to non-Bayesian agent behavior. (c) LLMs exhibit sophisticated strategic reasoning, highlighting both their potential and the societal risks they present.
@@ 2 Related Work
Our work builds on social learning, information design, online persuasion and RL, and LLMs.
@@
Social Learning Bikhchandani et al. (1992) and Banerjee (1992) introduced the sequential social learning framework in economics, and subsequent work has thoroughly characterized social-learning dynamics in more general settings (e.g., Smith and Sørensen (2000); Arieli and Mueller-Frank (2021)). This literature has also studied many variations in the kind of social information available to agents and various biases and misperceptions of the agents (see Dasaratha and He (2022); Bistritz et al. (2022); Eyster and Rabin (2010) as well as Bikhchandani et al. (2024) and the references therein).
@@
The literature on the control of social learning is relatively limited (Wei and Anastasopoulos, 2022; Smith et al., 2021; Krishnamurthy, 2012; Bhatt and Krishnamurthy, 2021). The closest works to our own are Wei and Anastasopoulos (2022) and Smith et al. (2021). Wei and Anastasopoulos (2022) assumes two-way communication between the planner and agents (i.e., agents can report their private signals to the planner in exchange for action recommendations). Smith et al. (2021) considers a benevolent planner who directly alters the choice rules of agents and also derives a decentralized pivot mechanism to improve learning. Our model does not rely on two-way communication between the agents and the planner and maintains the agents’ control of their actions, making it more suited for information-mediating algorithms which are often invisible or black-boxed to the users.
@@
Information Design In our setting, the planner optimizes their objective function by choosing the agents’ private signal precision. This is a constrained form of more classical information design questions (Bergemann and Morris, 2019). These models came into prominence with Kamenica and Gentzkow (2011)’s model of Bayesian persuasion — a planner chooses the signal structure informing a rational, Bayesian agent about an unknown state. Our planner’s problem is then a costly, constrained information design problem in each period, where the planner must choose a binary, symmetric signal at some cost. While many variations of information design have been studied, our primary distinction is to incorporate social learning between the agents.
There are very few works which consider information design problems in sequential social learning settings (Arieli et al., 2022; Wu et al., 2025). Both works consider a one-shot setting where the sender chooses an information structure at onset which is fixed for the sequence of agents. Arieli et al. (2022) considers a self-interested sender who chooses a general information structure as in the classic Bayesian Persuasion problem. Wu et al. (2025) constrain the information structure to be more informative in the sense of Blackwell and maximize the probability of a correct cascade. We consider a dynamic problem where the sender chooses a new structure for each agent in the sequence, engendering a more complex class of sender strategies.
Online Persuasion and RL There is growing recent work at the interface of online learning and Bayesian persuasion/information design. These works study online sequential persuasion under unknown environment parameters, in which a sender interacts repeatedly with short-lived receivers while lacking knowledge of the prior over states and/or receivers’ utility functions (e.g. Castiglioni et al. (2020); Bacchiocchi et al. (2024); Castiglioni et al. (2021); Agrawal et al. (2023)). A related set of papers examines sequential persuasion in dynamic environments or MDPs, introducing state transitions and/or foresight by the receiver; these models emphasize history-dependent or promise-based signaling policies and often connect to reinforcement learning (e.g. Wu et al. (2022); Su et al. (2022); Bernasconi et al. (2023)). Across these works, the focus lies in learning-based performance guarantees such as sublinear regret relative to the best signaling scheme in hindsight and/or approximation algorithms under computational constraints.
In contrast, our model does not require the planner to learn unknown parameters; sequential dependence arises from social learning among receivers, rather than from the sender updating beliefs about the environment. Moreover, the planner is not informationally advantaged: it does not observe the underlying state or private signals. Within this setting, we are able to characterize optimal policies for the planner. While sequential, our setting is not an online-learning or RL environment: actions do not modify the underlying payoff-relevant state (no reward externalities), and the planner does not learn unknown environmental parameters. The planner updates its knowledge only through observing agents’ actions, which reflect social learning and belief propagation rather than payoff exploration or state evolution. Sequential dependence therefore arises from information externalities, as agents’ actions shape the beliefs of future agents.
@@
LLMs More recently, some works have specifically studied LLMs in the context of information design (Harris et al., 2025; Cheng and You, 2025; Raifer et al., 2022; Bai et al., 2024; Duetting et al., 2025). These works all consider one-on-one interactions and, as such, do not capture the social dynamics we aim to investigate. The most similar work in this area is Duetting et al. (2025). They study information design where a sender chooses both a signal structure and a ’framing’ that conditions the posterior belief of the receiver, possibly non-Bayesian. They implement an LLM-based oracle approach for the computationally difficult task of optimizing over the framing space. This builds on the notion that LLMs can mimic humans as economic agents (Dillion et al., 2023; Horton, 2023). Their work does not address our main considerations of repeated interactions or social-learning; but we adapt parts of their methodology for our empirical study in Section 6.
@@ 3.1 Agents’ Decision Problems
Informed by $\mathcal{H}_{i}$ (equivalently, by the Markov property, $b_{i}$) and $q_{i}$, agent $i$ chooses action $a_{i}\in\Omega$ so as to maximize her utility. If both actions fetch the same utility, she chooses the action that matches $s_{i}$. Agent $i$ obtains a private belief $\tilde{b}_{i}$ about $\omega$ using $\mathcal{H}_{i}$ and her private signal $s_{i}$ of precision $q_{i}$ as follows (derivation in Appendix B.1):
$\tilde{b}_{i}$ $=\operatorname{\mathbb{P}}(\omega=G|\mathcal{H}_{i},q_{i},s_{i})=\begin{cases}\frac{q_{i}}{1+2b_{i}q_{i}-b_{i}q_{i}}b_{i}&s_{i}=G\\ \frac{1-q_{i}}{b_{i}+q_{i}-2b_{i}q_{i}}b_{i}&s_{i}=B\end{cases}$ (1)
Using this private belief, she chooses the action corresponding to the state of the world that is more likely as per her posterior belief $\tilde{b}_{i}$. In other words, she chooses $a_{i}=G$ if $\tilde{b}_{i}>0.5$; $a_{i}=B$ if $\tilde{b}_{i}<0.5$; and $a_{i}=s_{i}$ if $\tilde{b}_{i}=0.5$. Substituting for $\tilde{b}_{i}$, we find:
@@ 3.2 The Planner’s Problem
The optimal utility and policy of the altruistic planner are then defined as the supremum and arg supremum of $V_{A}^{\pi}(\cdot)$ over $\pi\in\Pi$ where $\Pi$ is the set of all policies. From this formulation, the myopic value and policy are recovered by setting $\delta=0$.
@@
The difference between the biased planner’s problem and the altruistic planner’s is in their objectives and, therefore, in the cost and reward functions. Refer to the example for a political campaign in Appendix A for elucidation of a biased planner. The biased planner seeks to induce action $G$ from each agent regardless of $\omega.$ When an agent chooses action $G$ (respectively, $B$), the planner, incurs cost 0 (respectively, $C>0$), regardless of $\omega$. The biased planner can make a private signal more or less precise than the baseline value of $p$, both of which incur costs.Any choice of precision other than $p$ incurs a cost for the biased planner as it requires him to tailor the ad to an agent, which in turn needs research on how the agent best understands any content. The biased planner incurs cost $\beta(|q_{i}-p|)$ for choosing signal precision $q_{i}$, with $\beta(0)=0$.
@@
There are a few noteworthy assumptions we make in our model: (1) The planner has the same history as the agents. This is limiting when the planner could have richer data, but not restrictive when all parties rely on the same public signals, such as early findings or survey data. (2) The planner can change agents’ signal precision, but signals are still passed through a binary symmetric channel. This is limiting if planners can cherry-pick, censor, or falsify signals, but not restrictive when their role is only to improve the quality of noisy data sources without altering content. (3) The planner’s control choices are fully observable. This is limiting when covert policies or framing are possible, but not restrictive in more transparent environments.
@@ 4 Optimal Altruistic Policies
The proof of Theorem 2 (Appendix C.3) is quite involved and may be of independent interest. The challenge is rooted in the dependence of agents’ actions on the public belief. In contrast, if the actions did not depend on the public belief process (e.g., as in Nyarko (1994)), the expected utility is linear function of the belief state, and the convexity of the value function then directly follows.
The convexity of the value function is instrumental in characterizing the optimal policy.
@@
Furthermore, if $t_{M}<0.5$, then $d_{A}<t_{M}$. See C.5 for the proof and figure 2(a) for an example.
@@ 5 Optimal Biased Policies
There exist $t_{1},\dots,t_{5}\in(0,p)$, $t_{1}<1-p\leq t_{2}\leq t_{3}<0.5\leq t_{4}\leq t_{5}<p$ so that:
@@
Theorem 5 (proven in Appendix C.8 and depicted in figure 2(a)) shows five possible phases.
@@ 6 Evaluation via LLM-based Simulation
We now utilize LLMs to empirically study the implications and applicability of our theoretical results and model. Our evaluation proceeds in three steps: we first analyze the behavior of LLM agents to identify key deviations from Bayesian rationality, then examine how an LLM planner adapts its strategy in response, and finally evaluate the resulting impact on social welfare.
To simulate the controlled social learning setting, we operationalize our model in a scenario where agents decide whether to buy a new car. As shown in figure 1(a), we employ LLMs in three roles. We describe each role below with corresponding prompts in Appendix E.2.2:
Agent: Role-playing as an assigned profile, agents observe previous actions and a private message about the car of known precision, form their belief about the car’s quality, and decide to buy or not.
Planner: After observing the history of actions, the planner selects the precision $q_{i}$ of agent $i$’s private signal according to their objective
Oracle: Given the choice of precision by the planner, an agent profile, and a fact sheet about the car (see Appendix E.2.1), the oracle generates a private signal of desired precision tailored to an agent.
Figure 1: (a) A system diagram for our experiments. The instance parameters detail are color-coded according to which of the LLM roles uses them. (b) Here we show the change in LLM (solid) and Bayesian (dashed) agents’ beliefs for a given prior after a positive (black) or negative (red) signal.
We compare the outcomes of these LLM-based simulations with numerical evaluations of the planner’s MDP. In both LLM and numerical experiments, we assume a linear cost function, $\beta(q)=k|q-p|$, and vary $k$, baseline precision $p$, and discount factor $\delta$, with the cost of an incorrect action $C$ fixed at $1$. See Appendix E for further detail on the experimental setup and prompting. In Appendix E.3, we validate both the beliefs and the performance of the oracle.
Our experiments reveal the following:
  1. 1.
    The importance and impact of social learning control: Planners (both analytical and LLM) can dramatically improve their own utility by accounting for and capitalizing on social learning dynamics. This may help or harm social welfare contingent upon the alignment between planner and agent objectives.
  2. 2.
    The robustness of our analytical characterization: LLM planners choose policies surprisingly similar to the non-obvious analytically optimal policies despite facing non-Bayesian agents in other LLMs.
  3. 3.
    The emergent strategic behavior of LLM: LLM planners exhibit sophisticated emergent strategic behavior in their choice of policies. Not only are LLM policies close to the optimal policies, they also seem to deviate in ways which account for the non-Bayesian nature of the LLM agents they face.